OpenStudy (anonymous):

What is the solution to the equation 2 ^(1 – 2x) = 3 ?

OpenStudy (lgbasallote):

change to log form first \[\implies \log_2 3 = 1 - 2x\] subtract 1 to both sides \[\implies \log_2 3 - 1 = -2x\] divide both sides by -2 \[\implies - \frac{\log_2 3 - 1}{2} = x\] does that help?

OpenStudy (anonymous):

IS IT -0.185?

OpenStudy (lgbasallote):

doesnt seem so...what did you do?